# On the complex zeros and the computational complexity of approximating the reliability polynomial

### v1

@@ 1. Case (1)
    Then by (5.2) $\omega_{1}:=x_{k}$ is the effective edge interaction of a two-terminal graph $G\in\mathcal{H}_{G_{0}}$, and satisfies $|\omega_{1}-\omega_{0}|\leq\varepsilon$. More precisely, letting $H_{0}=G_{0}^{\parallel n_{s}}$ and $H_{j}=(H_{j-1}\parallel G_{0})\bowtie G_{i_{j-1}}$, we have $G=H_{k}$. We can also compute its reliability polynomial, $R(G;p)$, using Lemma 2.1 in time $O(k)={\mathtt{poly}}({\mathtt{size}}(s,t,\varepsilon))={\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$. Additionally, $|G|=O(k)=O(\log(\varepsilon^{-1}))$ ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$.
@@ 2. Case (2)
    Take $n=\left\lceil\tfrac{|\omega_{0}-1|}{r}\right\rceil\in\mathbb{N}$ and note that $n\leq\tfrac{|p|+3}{r}+1$ and hence $n$ is constant in terms of the input. Then $u=\tfrac{\omega_{0}-1}{n}+1\in B(1,r)$. (Also, note that $u\in\mathbb{R}$ in case $p\in\mathbb{R}$.) We then run the algorithm from Case (1) with $s$ as defined above, $t=u$ and accuracy $\varepsilon/n$. The output $x$ is the effective edge interaction of some $H\in\mathcal{H}^{*}_{G_{0}}(p)$ and satisfies $|u-x|<\varepsilon/n$. The running time is bounded by ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon/n))={\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$ since $n$ is constant.
@@
    $|H^{\bowtie_{n}}|<n|H|={\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon)).$
    The running time is also bounded by ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$.
@@ 3. Case (3)
    By our pre-computation we know there exists $i\in\{1,\ldots,N\}$ and $x_{0}\in B(p,2)$ such that $g^{\circ i}(x_{0})=\omega_{0}$. We can compute it by computing the inverse $g^{-1}$ and determine which value $(g^{-1})^{\circ i}(\omega_{0})$ lies in $B(p,2)$. This takes only polynomial time in terms of ${\mathtt{size}}(\omega_{0})=O({\mathtt{size}}(y_{0}))$, since $N$ is constant. The idea is now to get an effective interaction $x_{1}$ that is close to $x_{0}$ using Case (2) and then apply $g^{\circ i}$ to it to obtain $\omega_{1}\in B(\omega_{0},\varepsilon)$. To make this precise, we first need to find out how close exactly we need to get to $x_{0}$.
@@
[…]
@@
    $|x_{0}-x_{1}|\leq\min\left\{\frac{\varepsilon^{\prime}|x_{0}-p|^{2}}{2|p-1|},|x_{0}-p|\right\}$
@@
    (5.4) ${\mathtt{size}}(\varepsilon^{\prime\prime})=O({\mathtt{size}}(\varepsilon,y_{0})).$
@@
    It follows from standard facts about algebraic numbers (cf. [Waldschmidtbook] and more specifically [main-approx]*Lemma 3.7) that the absolute logarithmic heights of $x_{0}-p$ and $\omega_{0}^{\prime}-1$ are bounded by $O({\mathtt{size}}(\omega_{0}))=O({\mathtt{size}}(y_{0}))$. Since $x_{0}\neq p$ and $\omega_{0}^{\prime}\neq 1$ it follows from [main-approx]*Lemma 3.7 that
    $\log(|x_{0}-p|)=O({\mathtt{size}}(y_{0}))\quad\text{and}\quad\log(|\omega_{0}^{\prime}-1|)=O({\mathtt{size}}(y_{0})).$
    This implies that ${\mathtt{size}}(\varepsilon^{\prime\prime})=O({\mathtt{size}}(\varepsilon,y_{0}))$.
    By the algorithm guaranteed from Case (2) we now find $x_{1}$ as the effective edge interaction of some two-terminal graph $H\in\mathcal{H}^{*}_{G_{0}}(p)$ such that $|x_{1}-x_{0}|\leq\varepsilon^{\prime\prime}$. Since ${\mathtt{size}}(x_{0})=O({\mathtt{size}}(y_{0}))$ the size of $H$ and the running time of the algorithm are both bounded by ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon^{\prime\prime}))$. By applying $g^{\circ i}$ to $x_{1}$ we find $\omega_{1}$ such that $|\omega_{0}-\omega_{1}|\leq\varepsilon$ and such that $\omega_{1}$ is the effective edge interaction of $G:=H\parallel G_{0}^{\parallel i}$. We can compute $R(G;p)$ using Lemma 2.1 in ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon^{\prime\prime}))$ time having access to $R(H;p)$ from the output of Case (2). Since ${\mathtt{size}}(\varepsilon^{\prime\prime})=O({\mathtt{size}}(\varepsilon,y_{0}))$. It follows that both the size of $H\parallel G_{0}^{\parallel i}$ and the running time of the algorithm are bounded by ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$, as desired.
    It remains to argue that $R(G;p)\neq 0$. Suppose towards contradiction that $R(G;p)=0$. Then, since $\omega_{1}=y_{G}(p)\neq 1$ (because $\varepsilon<r/2$ by our starting assumption in this case) it must be that $S(G;p)=0$. From this it follows by Lemma 2.1 that $S(H;p)=0$ since $S(G_{0};p)\neq 0$ by assumption. This implies that $x_{1}=y_{H}(p)=1$, as $R(H;p)\neq 0$ since $H\in\mathcal{H}^{*}_{G_{0}}(p)$. But then $\omega_{1}=g^{\circ i}(1)=1$, a contradiction.
@@ 5.4. Computing ratios
Let $p$ be an algebraic number. Suppose there exists an algorithm that on input of a (planar) graph $H$ computes a $0.25$-abs approximation (resp. $0.25$-arg approximation) to $R(H;P)$ in time polynomial in $|H|$.
Then there exists an algorithm that on input of a (planar) graph $H$ an edge $e$ of $H$, $G\in\mathcal{H}^{*}_{G_{0}}$ and the number $R(G;p)$ that computes a $0.25$-abs approximation (resp. $0.25$-arg approximation) to $R(H;p)+(y_{G}-(p+1))R(H\setminus e;)$ in time polynomial in $(|H|+|G|)$. In the planar setting we assume $G_{0}$ is planar with its terminals on the same face.
@@
  • there exists a ${\mathtt{poly}}(\log C,{\mathtt{size}}(y_{0},\varepsilon))$-time algorithm to compute, on input of $y_{0}\in\mathbb{Q}[i]$ and a rational number $\varepsilon>0$, an $0.25$-abs-approximation of $A\hat{y}+B$ for some algebraic number $\hat{y}\in B(y_{0},\varepsilon)$, or,
  • there exists a ${\mathtt{poly}}(\log C,{\mathtt{size}}(y_{0},\varepsilon))$-time algorithm to compute, on input of $y_{0}\in\mathbb{Q}[i]$ and a rational number $\varepsilon>0$, an $0.25$-arg-approximation of $A\hat{y}+B$ for some algebraic number $\hat{y}\in B(y_{0},\varepsilon)$.
Then, there exists an algorithm that, on input a rational $\delta>0$ and $C>0$ as above, outputs “A=0A=0” when $A=0$ and $B\neq 0$, and that outputs $A\neq 0$” and a number $y\in\mathbb{Q}[i]$, such that $B/A\in B_{\infty}(y,\delta/2)$ when $A\neq 0$. When $A=B=0$ it is allowed to output anything. The running time is ${\mathtt{poly}}(\log(C/\delta))$.
@@
    $\log(H\left(\frac{R(G;p)}{R(G\setminus e;p)}\right)\leq 2dm(\log(4)+h(p))$
@@
The proof of (b) follows along the same lines. We have by (a) and by [main-approx]*Lemma 3.7(b) $h(\frac{R(G;p)}{R(G\setminus e;p)})\leq 2m(\log(4)+h(p))-\log(4)-h(p)$. The claimed bound the follows from [main-approx]*Lemma 3.7(c) using that the ratio has degree at most $d$. ∎
@@
Let $C$ be the smallest integer bigger than $\exp(dm(\log(4)+h(p))$, where $m$ denotes the number of edges of $F$. Then by Lemma 5.7, both $|A|<C$ and $|B|<C$ as well as $|A|>1/C$ if $A\neq 0$ and $|B|>1/C$ if $B\neq 0$. We know show that Theorem 5.2 combined with Lemma 5.5 gives us the desired algorithm to be able to apply the box shrinking procedure (Theorem 5.6). Indeed on input of $y_{0}\in\mathbb{Q}[i]$ and $\varepsilon>0$ the algorithm of Theorem 5.2 gives us a two-terminal graph $G\in\mathcal{H}^{*}_{G_{0}}(p)$ such that $|y_{0}-(y_{G}-(p+1))|\leq\varepsilon$ in time ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$ as well as the value $R(G;p)$. Applying the algorithm of Lemma 5.5 we obtain an $0.25$-abs-approximation (resp. $0.25$-arg-approximation) to $R(F;p)+(y_{G}-(p+1))R(F-e;p)$ in time ${\mathtt{poly}}(|G|+|F|)={\mathtt{poly}}(\log(C),{\mathtt{size}}(y_{0},\varepsilon))$. We can therefore apply the box shrinking procedure (Theorem 5.6) with
$\delta=\exp(-(d^{2}+5d))C^{-2}.$
If the output of that algorithm is A=0"A"=0 we output, $b=0$ and r=1pr=1-p and if it outputs $A\neq 0"$ we output $b=1$ and we use a slight modification of the LLL [LLL]-based algorithm of Kannan, Lenstra and Lovász [kannan-1988] in the form of [main-approx]*Proposition 3.9. We input $d$, $H=2C$ and $y$ to that algorithm and in time ${\mathtt{poly}}(d,\log(C),{\mathtt{size}}(y))$ it outputs a polynomial $q$. We then output the algebraic number given by $q$ and the box $B_{\infty}(y,\delta/2)$.
@@ 5.5. Telescoping
Then there is an algorithm that on input of a (planar) graph $G$ computes $R(G;p)$ in ${\mathtt{poly}}(|G|)$ time.
@@
Let us first assume that $R(G;p)\neq 0$. We show by induction that $R(G_{i};p)\neq 0$ for all $i=0,\ldots,m$. Assuming $R(G_{i};p)\neq 0$ we know that not both $R(G_{i}-e_{i};p)=0$ and $R(G_{i}/e_{i};p)=0$. In case $R(G_{i}\setminus e_{i};p)\neq 0$ the algorithm outputs $b_{i}=1$ and hence $G_{i+1}=G_{i}-e_{i}$ and thus $R(G_{i+1};p)\neq 0$. In case $R(G_{i}-e_{i};p)=0$ and hence $R(G_{i}/e_{i};p)\neq 0$ the algorithm outputs $b_{i}=0$ and hence $G_{i+1}=G_{i}/e_{i}$ thus $R(G_{i+1};p)\neq 0$. Since $G_{m}$ has no edges, $R(G_{m};p)=c$ for all $p\in\mathbb{C}$, where $c=1$ if and only if $G_{m}$ has exactly one vertex, otherwise $c=0$. It thus follows that $c=1$. Since in case $b_{i}=0$ we have $R(G_{i}-e_{i};p)=0$ and thus by deletion contraction (Lemma 2.2) $R(G_{i};p)=(1-p)R(G_{i}/e_{i};p)=(1-p)R(G_{i+1};p)$, it follows that $r_{i}=\frac{R(G_{i};p)}{R(G_{i+1};p)}$ for all $i$. Therefore

### v2

@@ 1. Case (1)
    Then by (5.2) $\omega_{1}:=x_{k}$ is the effective edge interaction of a two-terminal graph $G\in\mathcal{H}_{G_{0}}$, and satisfies $|\omega_{1}-\omega_{0}|\leq\varepsilon$. More precisely, letting $H_{0}=G_{0}^{\parallel n_{s}}$ and $H_{j}=(H_{j-1}\parallel G_{0})\bowtie G_{i_{j-1}}$, we have $G=H_{k}$. We can also compute its reliability polynomial, $R(G;p)$, using Lemma 2.1 in time $O(k)={\mathsf{poly}}({\mathsf{size}}(s,t,\varepsilon))={\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$. Additionally, $|G|=O(k)=O(\log(\varepsilon^{-1}))$ ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$.
@@ 2. Case (2)
    Take $n=\left\lceil\tfrac{|\omega_{0}-1|}{r}\right\rceil\in\mathbb{N}$ and note that $n\leq\tfrac{|p|+3}{r}+1$ and hence $n$ is constant in terms of the input. Then $u=\tfrac{\omega_{0}-1}{n}+1\in B(1,r)$. (Also, note that $u\in\mathbb{R}$ in case $p\in\mathbb{R}$.) We then run the algorithm from Case (1) with $s$ as defined above, $t=u$ and accuracy $\varepsilon/n$. The output $x$ is the effective edge interaction of some $H\in\mathcal{H}^{*}_{G_{0}}(p)$ and satisfies $|u-x|<\varepsilon/n$. The running time is bounded by ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon/n))={\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$ since $n$ is constant.
@@
    $|H^{\bowtie_{n}}|<n|H|={\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon)).$
    The running time is also bounded by ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$.
@@ 3. Case (3)
    By our pre-computation we know there exists $i\in\{1,\ldots,N\}$ and $x_{0}\in B(p,2)$ such that $g^{\circ i}(x_{0})=\omega_{0}$. We can compute it by computing the inverse $g^{-1}$ and determine which value $(g^{-1})^{\circ i}(\omega_{0})$ lies in $B(p,2)$. This takes only polynomial time in terms of ${\mathsf{size}}(\omega_{0})=O({\mathsf{size}}(y_{0}))$, since $N$ is constant. The idea is now to get an effective interaction $x_{1}$ that is close to $x_{0}$ using Case (2) and then apply $g^{\circ i}$ to it to obtain $\omega_{1}\in B(\omega_{0},\varepsilon)$. To make this precise, we first need to find out how close exactly we need to get to $x_{0}$.
@@
    it follows that $|\omega_{0}-\omega_{1}|\leq\varepsilon$ (here we use $|\omega_{1}^{\prime}-1|\leq|\omega_{1}^{\prime}-\omega_{0}^{\prime}|+|\omega_{0}^{\prime}-1|$).
@@
    $|x_{0}-x_{1}|\leq\min\left\{\frac{\varepsilon^{\prime}|x_{0}-p|^{2}}{2|p-1|},|x_{0}-p|\right\},$
@@
    (5.4) ${\mathsf{size}}(\varepsilon^{\prime\prime})=O({\mathsf{size}}(\varepsilon,y_{0})).$
@@
    It follows from standard facts about algebraic numbers (cf. [Waldschmidtbook] and more specifically [main-approx]*Lemma 3.7) that the absolute logarithmic heights of $x_{0}-p$ and $\omega_{0}^{\prime}-1$ are bounded by $O({\mathsf{size}}(\omega_{0}))=O({\mathsf{size}}(y_{0}))$. Since $x_{0}\neq p$ and $\omega_{0}^{\prime}\neq 1$ it follows from [main-approx]*Lemma 3.7 that
    $\log(|x_{0}-p|)=O({\mathsf{size}}(y_{0}))\quad\text{and}\quad\log(|\omega_{0}^{\prime}-1|)=O({\mathsf{size}}(y_{0})).$
    This implies that ${\mathsf{size}}(\varepsilon^{\prime\prime})=O({\mathsf{size}}(\varepsilon,y_{0}))$.
    By the algorithm guaranteed from Case (2) we now find $x_{1}$ as the effective edge interaction of some two-terminal graph $H\in\mathcal{H}^{*}_{G_{0}}(p)$ such that $|x_{1}-x_{0}|\leq\varepsilon^{\prime\prime}$. Since ${\mathsf{size}}(x_{0})=O({\mathsf{size}}(y_{0}))$ the size of $H$ and the running time of the algorithm are both bounded by ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon^{\prime\prime}))$. By applying $g^{\circ i}$ to $x_{1}$ we find $\omega_{1}$ such that $|\omega_{0}-\omega_{1}|\leq\varepsilon$ and such that $\omega_{1}$ is the effective edge interaction of $G:=H\parallel G_{0}^{\parallel i}$. We can compute $R(G;p)$ using Lemma 2.1 in ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon^{\prime\prime}))$ time having access to $R(H;p)$ from the output of Case (2). Since ${\mathsf{size}}(\varepsilon^{\prime\prime})=O({\mathsf{size}}(\varepsilon,y_{0}))$. It follows that both the size of $H\parallel G_{0}^{\parallel i}$ and the running time of the algorithm are bounded by ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$, as desired.
    It remains to argue that $R(G;p)\neq 0$. Suppose towards contradiction that $R(G;p)=0$. Then, since $\omega_{1}=y_{G}(p)\neq\infty$, it must be that $S(G;p)=0$. From this it follows by Lemma 2.1 that $S(H;p)=0$ since $S(G_{0};p)\neq 0$ by assumption. This implies that $x_{1}=y_{H}(p)=1$, as $R(H;p)\neq 0$ since $H\in\mathcal{H}^{*}_{G_{0}}(p)$. But then $\omega_{1}=g^{\circ i}(1)=1$, contradicting our assumption that $1\notin B(\omega_{0},\varepsilon)$. This finishes the proof of Case (3).
@@ 5.4. Computing ratios
Let $p\neq 1$ be an algebraic number. Suppose there exists an algorithm that on input of a (planar) graph $H$ computes a $0.25$-abs approximation (resp. $0.25$-arg approximation) to $R(H;P)$ in time polynomial in $|H|$.
Then there exists an algorithm that on input of a (planar) graph $H$ and an edge $e$ of $H$, a two terminal graph $G\in\mathcal{H}^{*}_{G_{0}}(p)$, and the number $R(G;p)$ that computes a $0.25$-abs approximation (resp. $0.25$-arg approximation) to $R(H;p)+(y_{G}-(p+1))R(H\setminus e;)$ in time polynomial in $(|H|+|G|)$. In the planar setting we assume $G_{0}$ is planar with its terminals on the same face.
@@
  • there exists a ${\mathsf{poly}}(\log C,{\mathsf{size}}(y_{0},\varepsilon))$-time algorithm to compute, on input of $y_{0}\in\mathbb{Q}[i]$ and a rational number $\varepsilon>0$, an $0.25$-abs-approximation of $A\hat{y}+B$ for some algebraic number $\hat{y}\in B(y_{0},\varepsilon)$, or,
  • there exists a ${\mathsf{poly}}(\log C,{\mathsf{size}}(y_{0},\varepsilon))$-time algorithm to compute, on input of $y_{0}\in\mathbb{Q}[i]$ and a rational number $\varepsilon>0$, an $0.25$-arg-approximation of $A\hat{y}+B$ for some algebraic number $\hat{y}\in B(y_{0},\varepsilon)$.
Then, there exists an algorithm that, on input of a rational $\delta>0$ and $C>0$ as above, outputs “A=0A=0” when $A=0$ and $B\neq 0$, and that outputs $A\neq 0$” and a number $y\in\mathbb{Q}[i]$, such that $B/A\in B_{\infty}(y,\delta/2)$ when $A\neq 0$. When $A=B=0$ it is allowed to output anything. The running time and the size of $y$ are both ${\mathsf{poly}}(\log(C/\delta))$.
@@
    $\log(H\left(\frac{R(G;p)}{R(G\setminus e;p)}\right)\leq 2dm(\log(4)+h(p)).$
@@
The proof of (b) follows along the same lines. We have by (a) and by [main-approx]*Lemma 3.7(b), $h(\frac{R(G;p)}{R(G\setminus e;p)})\leq 2m(\log(4)+h(p))-\log(4)-h(p)$. The claimed bound then follows from [main-approx]*Lemma 3.7(c) using that the ratio has degree at most $d$. ∎
@@
Let $C$ be the smallest integer bigger than $\exp(dm(\log(4)+h(p))$, where $m$ denotes the number of edges of $F$. Then by Lemma 5.8, both $|A|<C$ and $|B|<C$ as well as $|A|>1/C$ if $A\neq 0$ and $|B|>1/C$ if $B\neq 0$. We know show that Theorem 5.2 combined with Lemma 5.6 gives us the desired algorithm to be able to apply the box shrinking procedure (Theorem 5.7). Indeed on input of $y_{0}\in\mathbb{Q}[i]$ and $\varepsilon>0$ the algorithm of Theorem 5.2 gives us a two-terminal graph $G\in\mathcal{H}^{*}_{G_{0}}(p)$ such that $|y_{0}-(y_{G}-(p+1))|\leq\varepsilon$ in time ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$ as well as the value $R(G;p)$. Applying the algorithm of Lemma 5.6 we obtain an $0.25$-abs-approximation (resp. $0.25$-arg-approximation) to $R(F;p)+(y_{G}-(p+1))R(F-e;p)$ in time ${\mathsf{poly}}(|G|+|F|)={\mathsf{poly}}(\log(C),{\mathsf{size}}(y_{0},\varepsilon))$. We can therefore apply the box shrinking procedure (Theorem 5.7) with
$\delta=\exp(-(d^{2}+5d))C^{-2},$
which has running time bounded by ${\mathsf{poly}}(\log(C/\delta))={\mathsf{poly}}(\log(C))$.
If the output of that algorithm is A=0"A"=0 we output, $b=0$ and r=1pr=1-p and if it outputs $A\neq 0"$ and $y\in\mathbb{Q}[i]$ we output $b=1$ and we use a slight modification of the LLL [LLL]-based algorithm of Kannan, Lenstra and Lovász [kannan-1988] in the form of [main-approx]*Proposition 3.9. We input $d$, $H=2C$ and $y$ to that algorithm and in time ${\mathsf{poly}}(d,\log(C),{\mathsf{size}}(y))={\mathsf{poly}}(\log(C))$ it outputs a polynomial $q$. We then output the algebraic number given by $q$ and the box $B_{\infty}(y,\delta/2)$. The overall running time is bounded by ${\mathsf{poly}}(\log(C))={\mathsf{poly}}({\mathsf{size}}(F))$.
@@ 5.5. Telescoping
Then there is an algorithm that on input of a (planar) graph $G$ computes $R(G;p)$ in ${\mathsf{poly}}(|G|)$ time.
@@
Let us first assume that $R(G;p)\neq 0$. We show by induction that $R(G_{i};p)\neq 0$ for all $i=0,\ldots,m$. Assuming $R(G_{i};p)\neq 0$ we know that not both $R(G_{i}-e_{i};p)=0$ and $R(G_{i}/e_{i};p)=0$. In case $R(G_{i}\setminus e_{i};p)\neq 0$ the algorithm outputs $b_{i}=1$ and hence $G_{i+1}=G_{i}-e_{i}$, and thus $R(G_{i+1};p)\neq 0$. In case $R(G_{i}-e_{i};p)=0$ and hence $R(G_{i}/e_{i};p)\neq 0$ the algorithm outputs $b_{i}=0$ and hence $G_{i+1}=G_{i}/e_{i}$, and thus also in this case $R(G_{i+1};p)\neq 0$. Since $G_{m}$ has no edges, $R(G_{m};p)=c$ for all $p\in\mathbb{C}$, where $c=1$ if and only if $G_{m}$ has exactly one vertex, otherwise $c=0$. It thus follows that $c=1$. Since in case $b_{i}=0$ we have $R(G_{i}-e_{i};p)=0$ and thus by deletion contraction (Lemma 2.2) $R(G_{i};p)=(1-p)R(G_{i}/e_{i};p)=(1-p)R(G_{i+1};p)$, it follows that $r_{i}=\frac{R(G_{i};p)}{R(G_{i+1};p)}$ for all $i$. Therefore